# Junzhou Huang8, Faez Ahmed9, Akash Srivastava2 & Dimitris N. Metaxas1 1

### v1

@@ Abstract
Discrete diffusion models have achieved success in tasks like image generation and masked language modeling but face limitations in controlled content editing. We introduce DICE (Discrete Inversion for Controllable Editing), the first approach to enable precise inversion for discrete diffusion models, including multinomial diffusion and masked generative models. By recording noise sequences and masking patterns during the reverse diffusion process, DICE enables accurate reconstruction and flexible editing of discrete data without the need for predefined masks or attention manipulation. We demonstrate the effectiveness of DICE across both image and text domains, evaluating it on models such as VQ-Diffusion, Paella, and RoBERTa. Our results show that DICE preserves high data fidelity while enhancing editing capabilities, offering new opportunities for fine-grained content manipulation in discrete spaces. Code is available at [link].
@@ 1 Introduction
Diffusion models have emerged as a powerful class of generative models, achieving remarkable success in high-fidelity image and video synthesis (Ho et al., 2020; Dhariwal & Nichol, 2021; Rombach et al., 2022; Ramesh et al., 2022; Ho et al., 2022; OpenAI, 2024). These models generate data by iteratively denoising samples from a simple noise distribution, effectively reversing a diffusion process that gradually corrupts data. Broadly, diffusion models can be categorized into continuous and discrete types, each tailored to different data modalities and applications.
Continuous diffusion models operate in continuous spaces, leveraging stochastic differential equations (SDEs) or their deterministic counterparts, ordinary differential equations (ODEs), to model the forward and reverse diffusion processes (Song et al., 2020; 2021). Advances such as flow matching (Lipman et al., 2022; Liu et al., 2022) have enhanced their efficiency and flexibility. These models have been successfully applied in various domains, including image editing (Meng et al., 2021; Avrahami et al., 2022; Mokady et al., 2022; Han et al., 2023; 2024; Zhang et al., 2023), medical imaging (He et al., 2023), and solving inverse problems (Chung et al., 2022; Stathopoulos et al., 2024). In image editing, continuous diffusion models enable controlled manipulation of images while preserving consistency with the underlying data distribution. A key capability enabling this is inversion—the process of reversing the diffusion model to recover the original noise vector or latent representation that could have generated a given data sample. Two main inversion approaches exist: deterministic inversion using ODEs (e.g., DDIM Inversion (Song et al., 2021)) and stochastic inversion by recording noise sequences (e.g., CycleDiffusion (Wu & De la Torre, 2022), DDPM Inversion (Dhariwal & Nichol, 2021)).
Discrete diffusion models are designed for inherently discrete data such as text or image tokens (Esser et al., 2021b). They adapt the diffusion framework to discrete spaces by defining appropriate transition kernels that corrupt and restore discrete data (Hoogeboom et al., 2021; Austin et al., 2021; Gu et al., 2022). Prominent examples include multinomial diffusion (Hoogeboom et al., 2021; Gu et al., 2022), D3PM (Austin et al., 2021), and masked generative models like MaskGIT (Chang et al., 2022), Muse (Chang et al., 2023). Despite their success in generation tasks, discrete diffusion models face limitations in controlled content editing. For instance, masked generative models achieve image editing through masked inpainting, where regions are masked and regenerated based on new conditions. However, this approach lacks the ability to inject information from the masked area into the inpainting process, limiting fine-grained control over the editing outcome (as illustrated in Figure 1).
@@
Figure 2: Here we demonstrate the two types of reconstruction and editing paradigms, namely ODE-based and Non-ODE based. (a,c) shows the ODE-based editing and reconstructions, while it provides accurate editing and reconstruction performances, it highly depends on the underlying ODE trajectory, which is not feasible in the discrete diffusion. However, the Non-ODE editing samples a trajectory by directly adding noise to $x_{0}$ and record the difference between the predicted $x_{t-1}$ and the sampled $x_{t-1}$ as indicated in the red arrow. In this way, we are able to reconstruct/edit the image without the strong condition of having an underlying ODE.
@@
Moreover, existing ODE-based inversion techniques developed for continuous diffusion models are not directly applicable to discrete diffusion models due to inherent differences in data representation and diffusion processes. This gap hinders the ability to perform precise inversion and controlled editing in discrete spaces. To address this challenge, we propose DICE (Discrete Inversion for Controllable Editing), the first inversion algorithm for discrete diffusion models to the best of our knowledge. Our method extends the stochastic inversion approach to discrete diffusion models, including both multinomial diffusion and masked generative models. The core idea is to record the noise sequence needed to recover a stochastic trajectory in the reverse diffusion process. Specifically, given an artificial trajectory where latent states have low correlation, we fit reverse sampling steps to this trajectory and save the residuals between targets and predictions. This process imprints the information of the original input data into the recorded residuals. During editing or inference, the recorded residuals are added back, allowing us to inject and control the amount of information introduced into the inference process.
@@
Our approach enables accurate reconstruction of the original input data and facilitates controlled editing without the need for predefined masks or attention map manipulation. It provides a flexible framework for fine-grained content manipulation in discrete spaces, overcoming the limitations of existing methods. We validate the effectiveness of DICE through extensive experiments on both image and text modalities. We evaluate our method on models such as VQ-Diffusion (Gu et al., 2022), Paella (Rampas et al., 2022), and RoBERTa (Liu et al., 2019), demonstrating its versatility across different types of discrete generative models. Additionally, we introduce a novel text-editing dataset to further showcase our method’s capabilities and to facilitate future research in this area. Our contributions can be summarized as follows:
@@
  • We introduce DICE, an inversion algorithm for discrete diffusion models, including multinomial diffusion and masked generative models. By recording and injecting noise sequences or masking patterns, DICE enables accurate reconstruction and controlled editing of discrete data without the need for predefined masks or attention manipulation.
@@ 2 Related Work
Discrete diffusion. D3PM (Austin et al., 2021) and Multinomial Diffusion (Hoogeboom et al., 2021) spearheaded the study of diffusion processes in discrete spaces by developing a corruption mechanism for categorical data. Following those works, Esser et al. (2021a) and Gu et al. (2022) introduced the VQ-GAN as a way to discretize the image into tokens. Also, extending to the natural language processing, Devlin et al. (2018) and Liu et al. (2019) proposed a bidirectional transformer for language understanding, which can be viewed as a discrete diffusion model (Wang & Cho, 2019). Additionally, Campbell et al. (2022) proposed discrete diffusion models with continuous time, while Lou et al. (2023) extended score matching (Song & Ermon, 2019) to discrete spaces by learning probability ratios. Gat et al. (2024) proposed discrete flow matching to extend the flow matching to discrete space. MaskGIT (Chang et al., 2022), Muse (Chang et al., 2023) and MMVID (Han et al., 2022) introduced efficient non-autoregressive methods for image generation by iteratively remasking and reprediction.
@@
Diffusion inversion. Diffusion inversion aims to find an encoding or latent representation of the input signal that can be used to reconstruct the original data. Traditional approaches to diffusion inversion are based on neural ODEs (Chen et al., 2018), such as DDIM inversion (Song et al., 2021) and flow matching (Lipman et al., 2022; Liu et al., 2022), where deterministic trajectories are used for inversion. Another class of methods focuses on stochastic differential equations (SDEs) (Song et al., 2020), including models like CycleDiffusion (Wu & De la Torre, 2022) and DDPM Inversion (Huberman-Spiegelglas et al., 2024), which rely on tracking noise or residuals along a stochastic path to recover the input. Our approach generalizes the concept of DDPM Inversion by extending it to discrete diffusion models, enabling effective inversion in both continuous and discrete settings.
@@
Inversion-based image editing. DDIM inversion (Song et al., 2021) has served as a foundational technique for various diffusion-based image editing approaches. In many image editing tasks, DDIM-type methods are often employed alongside guidance techniques like Prompt-to-Prompt (Hertz et al., 2022), which manipulate cross-attention maps. In contrast, DDPM inversion-based (Huberman-Spiegelglas et al., 2024) approaches are more user-friendly, as they do not require cross-attention manipulations. To address issues such as inaccurate reconstruction and error accumulation, Null-text Inversion (Mokady et al., 2022) introduces test-time optimization of null embeddings, ensuring the reconstruction trajectory aligns more closely with the DDIM inversion path. Negative-prompt Inversion (Miyake et al., 2023; Han et al., 2024) further improves time efficiency by providing a closed-form solution to an approximate inversion problem, reducing computational costs while maintaining competitive reconstruction quality.
@@ 3.1 Preliminaries
Masked generative modeling. Masked generative modeling is widely used in representation learning for both natural language processing and computer vision. It works by masking parts of the input and training the model to reconstruct the missing data. In models like BERT (Devlin et al., 2018) and RoBERTa (Liu et al., 2019), masked tokens ([MASK]) are predicted based on the surrounding context, excelling in text completion and embedding representation learning. For image generation, Paella (Rampas et al., 2022) adapts this approach for text-conditional image generation by renoising tokens instead of masking. The inference process in masked generative models typically involves iterative renoise/remask and repredict steps.
Multinomial Diffusion. Denoting $\bm{x}_{0}\in\{1,\ldots,K\}^{D}$ as a data point of dimension $D$. We use $\bm{v}({x}_{t}^{(i)})$ to denote the one hot column vector representation of the $i$-th entry of $\bm{x}_{t}$. To simplify notation, in the following we drop index $i$ and any function that operates on vector $\bm{x}_{t}$ is populated along its dimension. Diffusion model defines a markov chain $q(\bm{x}_{1:T}|\bm{x}_{0})=\Pi_{t=1}^{T}q(\bm{x}_{t}|\bm{x}_{t-1})$ that gradually add noise to the data $\bm{x}_{0}$ for $T$ times so that $\bm{x}_{T}$ contains little to no information. Discrete diffusion model (Hoogeboom et al., 2021; Austin et al., 2021; Gu et al., 2022) proposed an alternative likelihood-based model for categorical data, and defines the forward process following:
@@
with $p_{\theta}(\tilde{x}_{0}|x_{t})$ is parameterized by a neural network. We gradually denoise from $x_{T}$ to $x_{0}$ using 3. For numerical stability, the implementation uses log space instead of probability space. Masked generative models can be viewed as a special case of multinomial diffusion models with an additional absorbing state (or the [MASK] state). Its training objective can be viewed as a reweighted ELBO (Bond-Taylor et al., 2022).
@@ 3.2 Discrete Inversion for Controllable Editing
Non ODE-based inversion. ODE-based generative models, such as DDIM and flow matching, define an ODE trajectory. Due to the deterministic nature of ODEs, inversion can be achieved by solving the ODE using the Euler method in forward direction, ensuring reconstruction based on the inherent properties of the ODE. In contrast, another line of research focuses on SDE-based models, such as CycleDiffusion (Wu & De la Torre, 2022) and DDPM Inversion (Huberman-Spiegelglas et al., 2024). Broadly speaking, these approaches ensure reconstruction by recording the noises or residuals that are required to reproduce the stochastic trajectory. CycleDiffusion records the Gaussian noise $\bm{z}_{t}$ during sampling from posterior $p(\bm{x}_{t-1}|\bm{x}_{t},\bm{x}_{0}=\bm{x}_{0})$ and injects information of the input signal by feeding the true $\bm{x}_{0}$. DDPM Inversion, on the other hand, incorporates information into $\bm{z}_{t}$ by fitting the reverse process into an artificial stochastic trajectory obtained by independent q-sample. For both CycleDiffusion and DDPM Inversion, the key idea is to utilize the Gaussian reparameterization trick, $x=\mu+\sigma z\Leftrightarrow x\sim\mathcal{N}(x;\mu,\sigma^{2})$, and keeping track of the “noise” that could have generated the sample from mean. For discrete diffusion models, we utilize the Gumbel-Max trick (Maddison et al., 2014; Jang et al., 2016), $x=\operatorname*{arg\,max}{\left(\log(\bm{\pi})+\bm{g}\right)}\Leftrightarrow x\sim\text{Cat}(x;\bm{\pi})$. Figure 2 provides an intuition of the proposed method.
Inverting multinomial diffusion. Similar to Huberman-Spiegelglas et al. (2024), we start by sampling a stochastic trajectory, $\{\bm{x}_{t}\}$, a sequence of independent q-sample’s from $q(\bm{x}_{t}|\bm{x}_{0})$ (we populate the following sampling operation along the dimension of $\bm{x}_{t}$),
Note that here we use the Gumbel softmax trick (Jang et al., 2016), which is equivalent to sampling from categorical distribution $q(x_{t}|x_{0})$.
$\bm{y}_{t-1}=$ $\log(\text{onehot}(\bm{x}_{t-1})),~{}~{}\text{and}~{}~{}$
$\hat{\bm{y}}_{t-1}=$ $\log(\bm{\pi}_{\theta}(\bm{x}_{t},t)),$
$\bm{z}_{t}:=$ $\bm{y}_{t-1}-\hat{\bm{y}}_{t-1}$ (5)
Note that here the latent $\bm{z}_{t}\in\mathbb{R}^{D\times K}$.
@@
16:     $\bm{x}_{t-1}\leftarrow\tilde{\bm{x}}_{0}\odot(\bm{1}-\bm{m}_{t-1})+\bm{n}\odot\bm{m}_{t-1}$
@@
Inverting masked generative models. In masked generative modeling, the stochastic trajectory ${\bm{x}_{t}}$ is constructed according to the specific inference algorithm of the model in use. For example, in Paella Rampas et al. (2022), the masking is inclusive, meaning that as the time step $t$ increases, the set of masked tokens grows. In contrast, the Unleashing Transformer Bond-Taylor et al. (2022) employs random masking at each step, where masks are generated independently using the q-sample function. Without loss of generality, we define a denoiser function $\mathcal{D}_{\theta}$ (parameterized by $\theta$). This denoiser outputs the logits of the predicted unmasked data given the noisy tokens $\bm{x}_{t}$. Since in this case, the categorical sampling happens at sampling from the denoiser’s prediction, we therefore define an corresponding latent sequence:
@@
To provide more flexibility, we introduce the hyperparameters $\tau$, $\lambda_{1}$, and $\lambda_{2}$, which allow for finer control over the editing process. Specifically, $\tau$ represents the starting (and largest) timestep at which the editing process begins, while $\lambda_{1}$ controls the amount of information injected from the original input, and $\lambda_{2}$ governs the introduction of random noise.
Analysis. We describe a simple yet prototypical example of DDPM and compute the mutual information between encoded latents and the input signal.
Remark 3.1.
Given a simple Gaussian DDPM with ${\bm{x}}_{0}\sim\mathcal{N}({\bm{0}},{\bm{I}})$, latents $\{{\bm{z}}_{t}\}$ are obtained with DDPM inversion (Huberman-Spiegelglas et al., 2024), then the mutual information between ${\bm{z}}_{t}$ and ${\bm{x}}_{0}$ is:
The mutual information between $\bm{z}_{t}$ and $\bm{x}_{0}$ is illustrated in Figure 3. We observe that the amount of information encoded from $\bm{x}_{0}$ into $\bm{z}_{t}$ decreases as $t$ increases, motivating us to explore different scheduling strategies for $\lambda$’s (see Figure 7).

### v2

# Akash Srivastava2,3, Hongdong Li9, Junzhou Huang10, & Dimitris N. Metaxas1

@@ Abstract
Recent advances in discrete diffusion models have demonstrated strong performance in image generation and masked language modeling, yet they remain limited in their capacity for controlled content editing. We propose DICE (Discrete Inversion for Controllable Editing), a novel framework that pioneers precise inversion capabilities for discrete diffusion models, including both masked generative and multinomial diffusion variants. Our key innovation lies in capturing noise sequences and masking patterns during reverse diffusion process, enabling both accurate reconstruction and flexible editing without relying on predefined masks or attention-based manipulations. Through comprehensive experiments across image and text modalities using models such as Paella, VQ-Diffusion, and RoBERTa, we demonstrate that DICE successfully maintains high fidelity to the original data while significantly expanding editing capabilities. These results establish new possibilities for fine-grained content manipulation in discrete spaces.
@@ 1 Introduction
Diffusion models have emerged as a powerful class of generative models, achieving remarkable success in high-fidelity image and video synthesis [23, 12, 45, 42, 24, 39, 40, 10]. These models generate data by iteratively denoising samples from a simple noise distribution, effectively reversing a diffusion process that gradually corrupts data. Broadly, diffusion models can be categorized into continuous and discrete types, each tailored to different data modalities.
@@
Continuous diffusion models operate in continuous spaces, leveraging stochastic differential equations (SDEs) or their deterministic counterparts, ordinary differential equations (ODEs), to model the forward and reverse diffusion processes [49, 47]. Advances such as flow matching [29, 30, 9] have enhanced their efficiency and flexibility. These models have been successfully applied in various domains, including image editing [35, 3, 38, 18, 19, 59], medical imaging [20], and solving inverse problems [8, 50]. In image editing, continuous diffusion models enable controlled manipulation of images while preserving consistency with the underlying data distribution. A key capability enabling this is inversion—the process of reversing the diffusion model to recover the original noise vector or latent representation that could have generated a given data sample. Two main inversion approaches exist: deterministic inversion using ODEs (e.g., DDIM Inversion [47]) and stochastic inversion by recording noise sequences (e.g., CycleDiffusion [56], DDPM Inversion [12]).
Discrete diffusion models are designed for inherently discrete data such as text or image tokens [14]. They adapt the diffusion framework to discrete spaces by defining appropriate transition kernels that corrupt and restore discrete data [25, 2, 16]. Prominent examples include multinomial diffusion [25, 16], D3PM [2], and masked generative models like MaskGIT [5], Muse [6]. Despite their success in generation tasks, discrete diffusion models face limitations in controlled content editing. For instance, masked generative models achieve image editing through masked inpainting, where regions are masked and regenerated based on new conditions. However, this approach lacks the ability to inject information from the masked area into the inpainting process, limiting fine-grained control over the editing outcome, as illustrated in Figure 1.
@@
Figure 2: Here we demonstrate the two types of reconstruction and editing paradigms, namely ODE-based and Non-ODE based. (a,b) shows the ODE-based editing and reconstructions, while it provides accurate editing and reconstruction performances, it highly depends on the underlying ODE trajectory, which is not feasible in the discrete diffusion. However, the Non-ODE editing samples a trajectory by directly adding noise to $x_{0}$ and record the difference between the predicted $x_{t-1}$ and the sampled $x_{t-1}$ as indicated in the red arrow (c,d). In this way, we are able to reconstruct/edit the image without the strong condition of having an underlying ODE. (e,f) illustrate inversion and editing process for masked generative modeling (MGM) as in Algorithm 1.
@@
Moreover, existing ODE-based inversion techniques developed for continuous diffusion models are not directly applicable to discrete diffusion models due to inherent differences in data representation and diffusion processes. This gap hinders the ability to perform precise inversion and controlled editing in discrete spaces. To address this challenge, we propose DICE (Discrete Inversion for Controllable Editing), the first inversion algorithm for discrete diffusion models to the best of our knowledge. Our method extends the stochastic inversion approach to discrete diffusion models, including both multinomial diffusion and masked generative models. The core idea is to record the noise sequence needed to recover a stochastic trajectory in the reverse diffusion process. Specifically, given an artificial trajectory where latent states have low correlation, we fit reverse sampling steps to this trajectory and save the residuals between targets and predictions. This process imprints the information of the original input data into the recorded residuals. During editing or inference, the residuals are added back, allowing us to inject and control the amount of information introduced into the inference process.
@@
Our approach enables accurate reconstruction of the original input data and facilitates controlled editing without the need for predefined masks or attention map manipulation. It provides a flexible framework for fine-grained content manipulation in discrete spaces, overcoming the limitations of existing methods. We validate the effectiveness of DICE through extensive experiments on both image and text modalities. We evaluate our method on models such as VQ-Diffusion [16], Paella [43], and RoBERTa [31], demonstrating its versatility across different types of discrete generative models. Additionally, we introduce a novel text-editing dataset to further showcase our method’s capabilities and to facilitate future research in this area. Our contributions can be summarized as follows:
@@
  • We introduce DICE, an inversion algorithm for discrete diffusion models, including multinomial diffusion and masked generative models. By recording and injecting noise sequences or masking patterns, DICE enables accurate reconstruction and controlled editing of discrete data without predefined masks or attention manipulation.
@@ 2 Related Work
Discrete diffusion. Generative modeling over discrete spaces using diffusion models, analogous to continuous space diffusion, was first introduced in [46]. Argmax flows [25] and D3PM [2] further developed this concept by modeling the noising process as a discrete-time, discrete-space Markov chain. These approaches utilize a variational loss function to train a neural network to reverse the Markov chain. Following those works, [13] and [16] introduced the VQ-GAN as a way to discretize the image into tokens. This enables efficient non-autoregressive methods, including MaskGIT [5], Muse [6] and MMVID [17], for image generation by iteratively remasking and reprediction. Building on the continuous counterpart, which is trained via score matching as a regression task, recent work [32, 36] introduces a discrete analogue, termed ratio matching. This approach extends score matching [48] to discrete spaces by learning probability ratios. Continuing this trend, [15] proposed discrete flow matching. In natural language processing, BERT [11] and RoBERTa [31] can also be viewed through the lens of discrete diffusion [52], using a similar discrete sequence modeling approach.
@@
Diffusion inversion. Diffusion inversion aims to find an encoding or latent representation of the input signal that can be used to reconstruct the original data. Traditional approaches to diffusion inversion are based on neural ODEs [7], such as DDIM inversion [47] and flow matching [29, 30], where deterministic trajectories are used for inversion. Another class of methods focuses on stochastic differential equations (SDEs) [49], including models like CycleDiffusion [56] and DDPM Inversion [26], which rely on tracking noise or residuals along a stochastic path to recover the input. Our approach generalizes the concept of DDPM Inversion by extending it to discrete diffusion models, enabling effective inversion in both continuous and discrete settings.
@@
Inversion-based image editing. DDIM inversion [47] has served as a foundational technique for various diffusion-based image editing approaches. In many image editing tasks, DDIM-type methods are often employed alongside guidance techniques like Prompt-to-Prompt [21], which manipulate cross-attention maps, as well as self-attention maps, as demonstrated by approaches like Plug-and-Play [51], TF-ICON [33], and StyleAligned [22]. On the other hand, DDPM inversion-based approaches [26] are known for their user-friendly nature, as they typically do not require complex attention map manipulations. These approaches are also versatile and can integrate with semantic guidance techniques, such as SEGA and LEDITS++, enabling broader applicability. To address challenges such as inaccurate reconstruction and error accumulation, Null-text Inversion [38] introduces test-time optimization of null embeddings, ensuring the reconstruction trajectory aligns more closely with the DDIM inversion path. Negative-prompt Inversion [37, 19] further improves time efficiency by providing a closed-form solution to an approximate inversion problem, reducing computational costs while maintaining competitive reconstruction quality.
@@ 3.1 Preliminaries
Masked generative modeling. Masked generative modeling is widely used in representation learning for both natural language processing and computer vision. It works by masking parts of the input and training the model to reconstruct the missing data. In models like BERT [11] and RoBERTa [31], masked tokens ([MASK]) are predicted based on the surrounding context, excelling in text completion and embedding representation learning. For image generation, Paella [43] adapts this approach for text-conditional image generation by renoising tokens instead of masking. The inference process in masked generative models typically involves iterative renoise/remask and repredict steps.
Multinomial Diffusion. Denoting $\bm{x}_{0}\in\{1,\ldots,K\}^{D}$ as a data point of dimension $D$. We use $\bm{v}({x}_{t}^{(i)})$ to denote the one hot column vector representation of the $i$-th entry of $\bm{x}_{t}$. To simplify notation, in the following we drop index $i$ and any function that operates on vector $\bm{x}_{t}$ is populated along its dimension. Diffusion model defines a markov chain $q(\bm{x}_{1:T}|\bm{x}_{0})=\Pi_{t=1}^{T}q(\bm{x}_{t}|\bm{x}_{t-1})$ that gradually add noise to the data $\bm{x}_{0}$ for $T$ times so that $\bm{x}_{T}$ contains little to no information. Discrete diffusion model [25, 2, 16] proposed an alternative likelihood-based model for categorical data, and defines the forward process following:
@@
with $p_{\theta}(\tilde{x}_{0}|x_{t})$ is parameterized by a neural network. We gradually denoise from $x_{T}$ to $x_{0}$ using 3. For numerical stability, the implementation uses log space instead of probability space. Masked generative models can be viewed as a special case of multinomial diffusion models with an additional absorbing state (or the [MASK] state). Its training objective can be viewed as a reweighted ELBO [4].
@@ 3.2 Discrete Inversion for Controllable Editing
Non ODE-based inversion. ODE-based generative models, such as DDIM and flow matching, define an ODE trajectory. Due to the deterministic nature of ODEs, inversion can be achieved by solving the ODE using the Euler method in forward direction, ensuring reconstruction based on the inherent properties of the ODE. In contrast, another line of research focuses on SDE-based models, such as CycleDiffusion [56] and DDPM Inversion [26]. Broadly speaking, these approaches ensure reconstruction by recording the noises or residuals that are required to reproduce the stochastic trajectory. CycleDiffusion records the Gaussian noise $\bm{z}_{t}$ during sampling from posterior $p(\bm{x}_{t-1}|\bm{x}_{t},\bm{x}_{0}=\bm{x}_{0})$ and injects information of the input signal by feeding the true $\bm{x}_{0}$. DDPM Inversion, on the other hand, incorporates information into $\bm{z}_{t}$ by fitting the reverse process into an artificial stochastic trajectory obtained by independent q-sample. For both CycleDiffusion and DDPM Inversion, the key idea is to utilize the Gaussian reparameterization trick, $x=\mu+\sigma z\Leftrightarrow x\sim\mathcal{N}(x;\mu,\sigma^{2})$, and keeping track of the “noise” that could have generated the sample from mean. For discrete diffusion models, we utilize the Gumbel-Max trick [34, 27], $x=\operatorname*{arg\,max}{\left(\log(\bm{\pi})+\bm{g}\right)}\Leftrightarrow x\sim\text{Cat}(x;\bm{\pi})$. Figure 2 provides an intuition of the proposed method.
@@
16:     $\bm{x}_{t-1}\leftarrow\tilde{\bm{x}}_{0}\odot(\bm{1}-\bm{m}_{t-1})+\bm{n}\odot\bm{m}_{t-1}$triangleright Re-noise
@@
Method PSNR $\uparrow$ LPIPS×${}_{{}^{\times 10^{3}}}$downarrow MSE×${}_{{}^{\times 10^{4}}}$downarrow SSIM×${}_{{}^{\times 10^{2}}}$uparrow
Inpainting 10.50 565.11 1002.09 30.13
Ours 30.91 39.81 11.07 90.22
$\text{Ours}^{\dagger}$ Inf 0.07 0.01 99.99
Table 1: Inversion Reconstruction performance. $\dagger$ The metric is calculated between the original image and its inverted counterpart. Due to the encoding and decoding steps in the VQ-VAE/GAN process, some inaccuracies are introduced by the quantization. The PSNR is Inf due to the reconstruction of our method yielding the same VQ-VAE/GAN latents. Base model is Paella [43].
Inverting masked generative models. In masked generative modeling, the stochastic trajectory ${\bm{x}_{t}}$ is constructed according to the specific inference algorithm of the model in use. For example, in Paella [43], the masking is inclusive, meaning that as the time step $t$ increases, the set of masked tokens grows. In contrast, the Unleashing Transformer [4] employs random masking at each step, where masks are generated independently using the q-sample function. Without loss of generality, we define a denoiser function $\mathcal{D}_{\theta}$ (parameterized by $\theta$). This denoiser outputs the logits of the predicted unmasked data given the noisy tokens $\bm{x}_{t}$. Unlike DDPM or multinomial diffusion, where $x_{t-1}$ is not sampled from a posterior given $x_{t}$, the inference of masked modeling takes a different approach. In masked modeling, $x_{t}$ is obtained from sampled $\hat{x}_{0|t}$ by re-noising. Since the categorical sampling happens at sampling from the denoiser’s prediction, we therefore define an corresponding latent sequence:
@@
To provide more flexibility, we introduce the hyperparameters $\tau$, $\lambda_{1}$, and $\lambda_{2}$, which allow for finer control over the editing process. Specifically, $\tau$ represents the starting (and largest) timestep at which the editing process begins, while $\lambda_{1}$ controls the amount of information injected from the original input, and $\lambda_{2}$ governs the introduction of random noise (Algorithm 1 line 14).
Noise injection. We discuss three strategies as follows:
Linear. This is a natural form inspired by the Gumbel-Max trick: thinking of $\lambda_{1}\cdot\bm{z}$ as a correction term, then $\log(\bm{\pi})+\lambda_{1}\cdot\bm{z}$ is the corrected logit and $\lambda_{2}$ is the inverse of temperature of the logit to control the sharpness of the resulting categorical distribution, as
$\operatorname*{arg\,max}{(\log(\bm{\pi})+\lambda_{1}\cdot\bm{z}+\lambda_{2}\cdot\bm{g})}$
$=$ $\operatorname*{arg\,max}{(\frac{1}{\lambda_{2}}\left(\log(\bm{\pi})+\lambda_{1}\cdot\bm{z}\right)+\bm{g})},~{}~{}\lambda_{2}>0.$
$\text{Gumbel}(\mu_{G},\beta_{G}),\quad\text{with}$
$\beta_{G}$ $=\sqrt{\lambda_{1}^{2}\beta_{1}^{2}+\lambda_{2}^{2}\beta_{2}^{2}},$
where $\gamma\approx 0.58$ is the Euler-Mascheroni constant. We consider the variance preserving form:
Max. The third way is inspired by the property of Gumbel distribution [54], that if $G_{1}$, $G_{2}$ are iid random variables following $\text{Gumbel}(\mu,\beta)$ then $\max{\{G_{1},G_{2}\}}-\beta\log{2}$ follows the same distribution. We also consider the max function for noise injection:
We empirically find that linear strategy gives best results.
Inverting multinomial diffusion is more straightforward given its inference is similar to DDPM. We start by sampling a stochastic trajectory, $\{\bm{x}_{t}\}$, a sequence of independent q-sample’s from $q(\bm{x}_{t}|\bm{x}_{0})$ (we populate the following sampling operation along the dimension of $\bm{x}_{t}$),
Note that here we use the Gumbel max trick [27], which is equivalent to sampling from categorical distribution $q(x_{t}|x_{0})$.
$\bm{y}_{t-1}=$ $\log(\text{onehot}(\bm{x}_{t-1})),~{}~{}\text{and}~{}~{}$
$\hat{\bm{y}}_{t-1}=$ $\log(\bm{\pi}_{\theta}(\bm{x}_{t},t)),$
$\bm{z}_{t}:=$ $\bm{y}_{t-1}-\hat{\bm{y}}_{t-1}$ (6)
Note that here the latent $\bm{z}_{t}\in\mathbb{R}^{D\times K}$.
Analysis. We provide a quantitative analysis of mutual information in diffusion models using a simple, closed-form DDPM example. While not directly analyzing the discrete case, this study offers insight into how noise and latent injection affect information flow, motivating our scheduling strategy for $\lambda$ decay. Full details are provided in the supplementary materials.